ganeshie8 (ganeshie8):

parameterize the closed surface obtained by rotating the region inside the first quadrant enclosed by y = x^3 y = 4x about the x-axis

OpenStudy (kainui):

Find their two points of intersection. Imagine the axes laid out, or even better draw a picture so you can label x, r, and theta since this is just cylindrical coordinates. I guess show me where you're getting stuck since I know you're smart and I am probably just repeating stuff you know lol.

OpenStudy (perl):

i would use cylindrical coordinates as well, one moment

ganeshie8 (ganeshie8):

i learned how to deal with one surface but the two surfaces are confusing me at the moment

OpenStudy (perl):

lets use 'modified' cylindrical coordinates ( r , theta, x )

ganeshie8 (ganeshie8):

does that work when the surface is made up of two surfaces ?

OpenStudy (kainui):

Just do them separately @ganeshie8. One is the outside shell and one is simply the inside shell. Just find out when the two functions (radii) intersect.

ganeshie8 (ganeshie8):

My bad i have lost my latex work :(

ganeshie8 (ganeshie8):

they intersect at (0, 0) and (2, 8)

OpenStudy (kainui):

Yep, so you can just parametrize both in cylindrical coordinates then define your domain to be [0,2] and you're done, that's it.

ganeshie8 (ganeshie8):

x = u y = u^3cos(\theta) z = u^3sin(\theta) for rotating first curve i think

ganeshie8 (ganeshie8):

u = [0, 2]

ganeshie8 (ganeshie8):

we need a different parameterization for the other piece is it ? cant a single parameterization for the entire closed surface ?

OpenStudy (kainui):

Nope, not that I know of. Only a few nice geometric surfaces can be parameterized with a single function because their coordinate system is nice and matches up. For instance, the sphere requires a top and bottom in rectangular coordinates since the square root function is single valued.

OpenStudy (kainui):

You could no doubt define a new coordinate basis where it was possible to only define a single function, however it would be even more terrible than just gluing the two together like this

ganeshie8 (ganeshie8):

makes sense ! il try paameterizing the line, it shouldnt be hard i hope the same parameters \(r\) and \(\theta\) change smoothly across both the surfaces because this is a single closed surface.. im just worried splitting the parameterization breaks anything..

OpenStudy (perl):

this is a nice problem to go along with your last question on parametrization a surface of revolution

OpenStudy (perl):

parametrizing*

OpenStudy (kainui):

The derivative around the "seam" where the two connect together is not differentiable if that's what you mean @ganeshie8 It's just like a piecewise continuous function or differentiating absolute value.

ganeshie8 (ganeshie8):

thanks the two pieces are working perfectly !

OpenStudy (perl):

what was your final expression, can you post it ?

OpenStudy (perl):

i can try to graph it on my graphing calculator program , which can handle parametrically defined surfaces

ganeshie8 (ganeshie8):

perl i made two different parameterization as discussed above y=x^3 : ``` x = r y = r^3cos(\theta) z = r^3sin(\theta) ``` y=4x: ``` x = r y = 4rcos(\theta) z = 4rsin(\theta) ``` r : [0, 2] theta : [0, 2pi]

OpenStudy (perl):

thanks

OpenStudy (dan815):